Detailed topic list for PartnerObserving Web

Results from PartnerObserving web retrieved at 02:35 (GMT)

01Mar2012_Log
Observing Log for PISCES night of 01 March 2012 UT AZ / OSURC mini queue with Hill et al., Close et al., McGreer et al., Woodward et al., Observer: JHill (LBTO w...
01Mar2015_Log
Mar 1 (UT) The dome was never opened due to a bad weather. Comments
02Mar2012_Log
Observing Log for PISCES night of 02 March 2012 UT AZ / OSURC mini queue with Hill et al., Close et al., McGreer et al., Woodward et al., Observer: JHill (LBTO w...
03Dec2015_Log
Dec 3 (UT) Observers: Wen fai Fong PI programs: Fong 00:00 UT Dome opened. 00:37 UT Slew to sky flat field. Sky flat test. 00:43 UT Start sky flats. 01:08 U...
03Mar2012_Log
Observing Log for PISCES night of 03 March 2012 UT AZ / OSURC mini queue with Hill et al., Close et al., McGreer et al., Woodward et al., Observer: JHill (LBTO w...
03Mar2013_Log
Some clouds around, mostly clear. High frequency readout noise can be seen in the images. All darks before 0019 are bad blank mask was not in, light in dome goo...
03Oct2007
PartnerObserving Log for 2007 Oct 03 04 Observers: J. Hill, V. Testa Telescope Operator: A. Ceranski Never opened because of high humidity and eventually fog. T=...
r3 - 21 Jul 2009 - 22:55 by NormCushing
04Dec2015_Log
Dec 4 (UT) Observers: Wen fai Fong PI programs: Fong 00:00 UT Dome opened. 01:00 UT Focusing, stars have a large coma. Global offsets were causing comas. Lots...
04Mar2013_Log
Sky is completely clouded over, doesn't look to be getting better any time soon. Pickup noise in images from yesterday seems to have gone away. 3:25pm Start takin...
04Oct2007
PartnerObserving Log for 2007 Oct 04 05 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski This is scheduled SDT time whi...
r5 - 21 Jul 2009 - 22:55 by NormCushing
05Mar2013_Log
Observers: Walth, Penner, Clement mostly clear, seeing ~1.5" (guide star) 19:10 telluric star, GCS=1.35" 20:07 done with telluric for Greg's masks, moving to Kyl...
05Nov2007
PartnerObserving Log for 2007 Oct 06 07 Observers: J. Hill, A. Rakich Telescope Operator: D. Gonzalez Huerta (all the technical images aren't included in this log...
r2 - 21 Jul 2009 - 22:55 by NormCushing
05Oct2007
PartnerObserving Log for 2007 Oct 05 06 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski First images in 2007106 are th...
r4 - 21 Jul 2009 - 22:55 by NormCushing
06Mar2013_Log
18:18 overcast, taking darks and flats for last night observations. 22:15 conditions slightly improving, moving to GOODS N 22:30 seeing probably ok (guider ~1.2")...
06Oct2007
PartnerObserving Log for 2007 Oct 06 07 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski 005954 010215 5 bias frame...
r7 - 21 Jul 2009 - 22:55 by NormCushing
07Apr2013_Log
Log 2013 04 08 2:00 some clouds coming in at sunset from west 2:31 heading to Feige 34. seeing 1.2''. 2:40 observing Feige 34 with 5'' slit. clouds rolling in. 2:...
07Mar2013_Log
sky is clear, forecasts are good for tonight. 17:15 starting darks 19:04 we are open, moving to Kyle's flux std 19:21 collimated. guider avg~0.8". Starting acq. ...
07Oct2007
PartnerObserving Log for 2007 Oct 07 08 Observers: J. Hill, V. Testa Instrument Support: R. Bertram Telescope Operator: A. Ceranski 00:49 25 bias frames 01:05 ...
r7 - 21 Jul 2009 - 22:55 by NormCushing
08Apr2013_Log
Did not open due to high winds and oncoming storm. Main.IanMcGreer 10 Apr 2013
08Jun2007
PartnerObserving Log for 2007 June 08/09 Technical Observer: J. Hill Telescope Operator: D. Gonzalez Instrument: LBC Blue Technical observing with a bit of SDT M8...
r7 - 21 Jul 2009 - 22:55 by NormCushing
08Oct2007
PartnerObserving Log for 2007 Oct 08 09 Observers: J. Hill, V. Testa Telescope Operator: A. Ceranski Telescope and Instrument Support (alseep on site) 051 25 ...
r6 - 21 Jul 2009 - 22:55 by NormCushing
09Apr2013_Log
Did not open due to ice on dome after day long snow (just a few inches on the ground). Main.IanMcGreer 10 Apr 2013
09Jun2007
PartnerObserving Log for 2007 June 09/10 Technical Observer: J. Hill Telescope Operator: D. Gonzalez Instrument: LBC Blue Technical observing with a bit of SDT M8...
r8 - 21 Jul 2009 - 22:55 by NormCushing
09May2007
PartnerObserving Log for 2007 May 09/10 Technical Observers: J. Hill, E. Diolaiti, A. DiPaola, R. Speziali Telescope Operator: E. Lundin Technical Support: J. Lit...
r6 - 21 Jul 2009 - 22:55 by NormCushing
09Oct2007
PartnerObserving Log for 2007 Oct 09 10 Observers: V. Testa, J. Hill, Stephane Herbert Fort Telescope Operator: David Gonzalez Huerta Telescope and Instrument S...
r4 - 21 Jul 2009 - 22:55 by NormCushing
10Apr2007
PartnerObserving Log for 2007 Apr 10/11 Observers: John Hill, Mark Wagner Telescope Operator: David Gonzales Instrument: LBC Blue Some more SDT related observin...
r4 - 21 Jul 2009 - 22:55 by NormCushing
10Apr2013_Log
0150 UT\xA0\xA0\xA0\xA0 Sunset. Unable to open due to high humidity. 0209 UT\xA0\xA0\xA0\xA0 Beginning calibration sequence (bias8k) 0230 UT\xA0\xA0\xA0\...
10Apr2014_Log
April 9, 2014 19:36 Telluric acquisition, HIP61534 19:50 HIP61534 HK spec 20:09 CLJ1226 acquisition 20:21 CLJ1226 start HK spec, 1.0" seeing, non photometric,...
10Dec2012_Log
Main.Zheng 11 Dec 2012 6:45pm, Acqusition UDS mask. 7:10pm, Telescope has problem of tracking, and get fixed in 20min. 7:45pm, Start Science exposure, seein...
10Jan2011_Log
January 10, 2011 OBSERVERS: Jill Bechtold, David Sand, Marco Pedani, Michael Rutkowski, Russell Ryan OPERATOR: David Gonzales Huerta Afternoon: Some high clouds....
10Jun2007
PartnerObserving Log for 2007 June 10/11 Observers: V. Burwitz, J. Hill Telescope Operator: D. Gonzalez Instrument: LBC Blue Open after sunset. T=10.9degC, D= 3....
r6 - 21 Jul 2009 - 22:55 by NormCushing
10June2014_Log
June 10 * 20:13 (before twilight ends) start taking acquisition image on M1115 (long slit, zJ), clear, seeing 0.7" * 20:33 starting science on M1115 * 21:05 done ...
10Mar2007
PartnerObserving Log for 2007 Mar 10/11 Observer: John Hill Telescope Operator: Ed Lundin Some more SDT related observing during engineering time. Note that I'...
r7 - 21 Jul 2009 - 22:55 by NormCushing
10Mar2011_Log
Observers: Marie Rex, Wiphu Rujopakarn, Dave Thompson Clear, seeing .4 .5" 7:30 pm (2:30 UT) start 2 hours of H band science observations of MACS 0717 (Rex) Stil...
10Mar2016_Log
Mar 10 (UT) Observers: Dennis Zaritsky 01:15 UT Dome opened. all dual model, long slit, grating spectroscopy G191B2B standard sequence as spectrophotometric tar...
10May2007
PartnerObserving Log for 2007 May 10/11 Technical Observers: J. Hill, A. DiPaola, R. Speziali Telescope Operator: E. Lundin Instrument: LBC Blue Some more SDT rel...
r7 - 21 Jul 2009 - 22:55 by NormCushing
10Nov2015_Log
Nov 10 (UT) Observer: Wen fai Fong (PI: Smith)/Ramesh and Peter (PI: Stark) 00:36 UT Dome opened, sky flat tests. 01:29 UT Sky flats done, slew to first targe...
10Oct2007
Observers are: V. Testa, S. Herbert Fort, D. Gonzales Huerta Mostly sunny with few thin clouds in the afternoon 00:35 25 bias 003632 005038 01:00 Sunset Stil...
11-12February2007
PartnerObserving Log for 2007 Feb 11/12 Observer: Tom Herbst, Giacomo Beccari Inst. Support: John Hill TOs: Ed Lundin David Gonzalez Winter storm started around...
r3 - 21 Jul 2009 - 22:55 by NormCushing
11Apr2007
Observers: Jason Harris, Paul Martini, Jose Prieto TO: David Gonzalez Instrument Support: John Hill Almanac: Sunset ( 3000 m horizon): 18 55 MST; Sunrise: 5 45 ...
11Apr2013_Log
0010 UT Trying to wake MODS, but it hands during 'blue tedpower on' 0020 UT Attempting to reset MODS to clear the problem w/ 'tedpower' 0030 UT Test preset ...
11Apr2014_Log
April 10, 2014 19:15 Pointing and collimation check 19:20 Telluric acquisition HIP35050 19:30 Struggling to find correct star because no finder were provided 19:5...
11Dec2012_Log
Main. Dan, Zheng 11 Dec 2012 6:25pm, start acquisition on Fuyan's lens, seeing is ~0.9'', earlier condition is not suitable for twilight flat. 6:30pm, start ...
11June2014_Log
June 11 * 20:50 Cloudy, not open LUCI is still broken, switching to MODS * 21:15 Less clouds, we are opening * 21:32 start acquisition on CLJ1226m1, seeing 1" * 2...
11Mar2007
PartnerObserving Log for 2007 Mar 11/12 Observer: John Hill TO: Ed Lundin Some more bonus SDT observations in engineering time. As with last night, I'm only usin...
r4 - 21 Jul 2009 - 22:55 by NormCushing
11Mar2011_Log
Observers: Marie Rex, Wiphu Rujopakarn, Dave Thompson Too cloudy for twilight sky flats. 0.6" seeing. Taking test images with Br gamma filter detector looks sa...
11Mar2016_Log
Mar 11 (UT) Observers: Dennis Zaritsky XX:XX UT Dome opened. XX:XX UT Dome closed. Comments
11May2007
PartnerObserving Log for 2007 May 11/12 Technical Observers: J. Hill, A. DiPaola, R. Speziali Telescope Operator: E. Lundin Instrument: LBC Blue After waiting for...
r5 - 21 Jul 2009 - 22:55 by NormCushing
11Nov2015_Log
Nov 11 (UT) Observer: Ian McGreer 00:35 UT Sunset, buried in fog. Humidity 97%, wind gusting. Comments
11Oct2007
Observers: V. Testa, S. Herbert Fort, D. Gonzales Huerta Quite nice and sunny in the afternoon. Sky looks very clear.At the moment it is candidate photometric. 06...
Number of topics: 50
Page 1 of 30 Next >

See also the faster WebTopicList
Topic revision: r1 - 26 Feb 2018, ProjectContributor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback